OpenStudy (anonymous):

DUE IN 10 MINUTES, PLEASE HELP ME!!!!!!! using only the digits 1-8 how many 6 digit numbers can be formed if adjacent digits must be different

OpenStudy (anonymous):

i multiplied 8*7*7*7*7*7 and got 134,156

OpenStudy (cwrw238):

i think its 8*7*6*5*4*3

OpenStudy (anonymous):

nope because you can repeat numbers

OpenStudy (anonymous):

as long as they are not adjacent

OpenStudy (cwrw238):

its beenn such a long time

OpenStudy (cwrw238):

oh i see - ok that wrong then

OpenStudy (anonymous):

whole number? :D lol, this is a crazy problem

OpenStudy (anonymous):

Its not what cwrw said??

OpenStudy (anonymous):

no because you can repeat numbers

OpenStudy (anonymous):

O_o... Uh..

ganeshie8 (ganeshie8):

@KingGeorge

OpenStudy (anonymous):

Ok, I got it. Add to the list if am missing anything... 8^3 * 7^3 = 175616 2 combo 8^3 * 7^2 * 6 = 150528 6 Combo 8^3 * 7^2 * 5 = 125440 6 Combo 8^3 * 7^2 * 4 = 100352 6 Combo 8^3 * 6^2 * 7 = 129024 6 Combo 8^3 * 7 * 6 * 5 = 107520 (this is the hardest one) 8^2 * 7^3 * 6 = 131712 6 Combo 8^2 * 7^3 * 5 = 109760 6 Combo

OpenStudy (anonymous):

8^3 * 7 * 6 * 5 = 107520 12 Combination total Mirror 8*7*8*6*8*5 5*8*6*8*7*8 8*7*8*5*8*6 ... 8*6*8*7*8*6 8*6*8*6*8*7 8*5*8*6*8*7 8*5*8*7*8*6 Sum of 12 8*_*8*_*_*8 8*_*_8_*8 Looks like 12 comb total of 24 for this pattern 62 in totals...?

OpenStudy (anonymous):

thats not correct because of the fact the you would have adjacent numbers next to each other

OpenStudy (anonymous):

adjacent number like...? lol

OpenStudy (raden):

it like ababab. so, it would be 8 * 7 * 7 * 7 * 7 * 7

OpenStudy (anonymous):

:? I was under the assumption that 8787878 is ok since the "adjacent number is different "87..."

OpenStudy (raden):

8 * 7 * 7 * 7 * 7 * 7 = ... that was my calculation

OpenStudy (raden):

the first digit, obvious we can select 8 numbers 2nd digit we only can select 7 numbers (so that no dupication here) 3th digit, we can select 7 numbers (remember there are 8 numbers, so if the left number is 1, we can choce 2,3,4,5,6,7,8) 4th digit, still we can select 7 numbers and so on ....